OpenStudy (anonymous):

In how many ways can a committee of size 3 be formed from 5 people? Assume that, when people are chosen for a committee, the order of the choices does not matter.

OpenStudy (btaylor):

using combinations: \[choices = _{5}C_{3} = 10 ways\]

OpenStudy (anonymous):

thanks is their a way to do in calculator?

OpenStudy (anonymous):

no need \[\frac{5\times 4}{2}=5\times 2=10\]

OpenStudy (btaylor):

@CaseyC there is, depending on your calculator. What type do you have?

OpenStudy (anonymous):

but yes, most calculators have this feature. depends on the calculator though

OpenStudy (anonymous):

In general \[\binom n r =\frac{ n\times(n-1)\times(n-2)\times(n-3)\times\cdots \times(n-r+1)}{r!} \]

OpenStudy (anonymous):

14,190?

OpenStudy (zarkon):

14,190...wow

OpenStudy (anonymous):

that is my kind of mistake!

OpenStudy (btaylor):

@vic1 Casey C wants combinations, not permutations.